OpenStudy (anonymous):

i will give metal and be fan...solve for m? 3m/7=21

OpenStudy (anonymous):

3m = 7 * 21

OpenStudy (anonymous):

147???

OpenStudy (anonymous):

but 3 xm=21 do you think it would be 7 /21?

OpenStudy (anonymous):

M=49

OpenStudy (anonymous):

can you show me how it would be 49 please?

OpenStudy (anonymous):

7*21=147 then 147/3=49

OpenStudy (anonymous):

ohh... but how does that equal 21? can you please show me?:)

OpenStudy (anonymous):

3x49/7=21

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

welcome.